OpenStudy (anonymous):

Find the zeros of EACH function. F(x)=x^2-12x+45 And g(x)=x^2+6x+34

OpenStudy (anonymous):

For both, you have to factor the quadratic equations into the individual binomials (x + a) * (x + b) so that a * b is equal to the last coefficient and a + b is equal to the middle coefficient. e.g. (x + 1) (x + 2) = x^2 + 3x + 2 where a = 1, b = 2 After that, you simply solve it while making it equal to zero. e.g. (x + 1) (x + 2) = 0 x = -1, -2. If you can't figure out the coefficients, use the quadratic formula, x = (-b +/- sqrt(b^2 - 4ac)) / 2a. This takes more work, but gives you the zeroes directly.